OpenStudy (anonymous):

x^2-8x+16=0 Can someone show me step-by-step how to solve this equation?

OpenStudy (anonymous):

e^ln(x^2-8x+16)=e^0 x^2-8x+16=1 x^2-8x+15=0 (x-3)(x-5)=0 x=3,5

OpenStudy (anonymous):

Could you just explain what you did there?

OpenStudy (anonymous):

Raise both with base e so you get: x^2-8x+16 =e^0 x^2 - 8x + 16 = 1 x^2 - 8x + 15 = 0 (x-5) = 0 (x-3) = 0 x = 5 x = 3

OpenStudy (anonymous):

You have a perfect square. You are looking for 2 numbers that multiply to +16 and add to -8. Those 2 numbers are -4 and -4 (-4 twice). This fits the form of: (x - 4)(x - 4) or (x - 4)^2 = 0 NOT 5 and 3.

OpenStudy (anonymous):

There really is only one step. It consists of recognizing the completion of a perfect square once one looks at: x^2 - 8x At that point, we take half of -8 and square that to get 16. But 16 is what we already have left to complete that square. You could do: (x^2 - 8x + 16) + 16 = 0 + 16 and then subtract 16 from both sides, but that step is not necessary.